OpenStudy (yusufm):

how do you prove 2/(sqt3cosx+sinx)=sec(pi/6-3)

OpenStudy (yusufm):

\[2/(\sqrt{3}cosx+sinx)=\sec(\pi/6-3)\]

OpenStudy (zzr0ck3r):

that would imply the thing on the left is a constant...